# Constructing an Optimal Behavior Basis for the Option Keyboard

### v1

@@ 1 Introduction
Reinforcement learning (RL) methods have been successfully used to solve complex sequential decision-making problems (Silver et al., , 2017; Bellemare et al., , 2020). However, traditional RL algorithms typically require thousands or millions of interactions with the environment to learn a single policy for a single task. Multi-task RL methods address this limitation by enabling agents to quickly identify solutions for new tasks with minimal or no additional interactions. Such methods often achieve this by learning a set of specialized policies (or behavior basis) designed for specific tasks and subsequently combining them to more rapidly solve novel tasks.
@@
A powerful approach for combining policies to solve new tasks in a zero-shot manner leverages successor features (SFs) and generalized policy improvement (GPI) (Barreto et al., , 2017, 2018, 2020). GPI can combine base policies to solve a new task, producing a policy that is at least as good as any individual base policy. Importantly, however, GPI policies are not guaranteed to be optimal. Optimality can be ensured if tasks can be expressed as a linear combination of reward features using techniques that compute a convex coverage set (CCS; Alegre et al., , 2022). A CCS is a set of policies that includes an optimal policy for any linear task. Unfortunately, the number of policies in a CCS often grows exponentially with the number of reward features (Roijers, , 2016), making existing algorithms computationally expensive and difficult to scale to complex domains (Yang et al., , 2019; Alegre et al., , 2022; Alegre et al., 2023b, ).
@@
The option keyboard (OK) method improves upon GPI by producing policies that are at least as good—and often better. It achieves this through a learned meta-policy that dynamically combines base policies by assigning state-dependent linear weights to reward features (Barreto et al., , 2019). This allows OK to express a larger spectrum of behaviors than GPI and potentially solve tasks beyond linear rewards, enabling it to produce policies that GPI cannot represent. However, its performance critically depends on the choice of base policies available. This raises a key question and open problem: is there an optimal set of base policies for the OK—an optimal behavior basis—that enables zero-shot identification of optimal solutions for any linear tasks? Existing works using OK assume that a set of base policies is either known a priori or constructed with expert domain knowledge and do not address the problem of identifying a good behavior basis (Barreto et al., , 2019; Carvalho et al., 2023b, ).
@@
We solve this open problem by introducing Option Keyboard Basis (OKB), a novel method with strong formal guarantees that efficiently identifies an optimal behavior basis for the OK. Importantly, our method provably identifies a set of policies that allows the OK to optimally solve any linear task; i.e., it can express all policies in a CCS. We show that it significantly reduces the number of base policies required to ensure zero-shot optimality in new tasks. Furthermore, we prove that the set of policies it can express is strictly more expressive than a CCS, enabling particular classes of non-linear tasks to be solved optimally. We empirically evaluate our method in challenging high-dimensional RL problems and show that it consistently outperforms state-of-the-art GPI-based approaches. Importantly, we also observe that the performance gain over competing methods becomes more pronounced as the number of reward features increases.
@@ 2.1 Reinforcement Learning
An RL problem (Sutton and Barto, , 2018) is typically modeled as a Markov decision process (MDP). An MDP is defined as a tuple $M\triangleq(\mathcal{S},\mathcal{A},p,r,\mu,\gamma)$, where $\mathcal{S}$ is a state space, $\mathcal{A}$ is an action space, $p(\cdot|s,a)$ describes the distribution over next states given that the agent executed action $a$ in state $s$, $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function, $\mu$ is an initial state distribution, and $\gamma\in[0,1)$ is a discounting factor. Let $S_{t}$, $A_{t}$, and $R_{t}\triangleq r(S_{t},A_{t},S_{t+1})$ be random variables corresponding to the state, action, and reward, respectively, at time step $t$. The goal of an RL agent is to learn a policy $\pi:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted sum of rewards (return), $G_{t}=\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}$. The action-value function of a policy $\pi$ is defined as $q^{\pi}(s,a)\triangleq\mathbb{E}_{\pi}[G_{t}|S_{t}=s,A_{t}=a]$, where $\mathbb{E}_{\pi}[\cdot]$ denotes the expectation over trajectories induced by $\pi$. Given $q^{\pi}$, one can define a greedy policy $\pi^{\prime}(s)\in\operatorname*{arg\,max}_{a}q^{\pi}(s,a)$. It is guaranteed that $q^{\pi^{\prime}}(s,a)\geq q^{\pi}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}$. The processes of computing $q^{\pi}$ and $\pi^{\prime}$ are known, respectively, as the policy evaluation and policy improvement steps. Under certain conditions, repeatedly executing the policy evaluating and improvement steps leads to an optimal policy $\pi^{*}(s)\in\operatorname*{arg\,max}\nolimits_{a}q^{*}(s,a)$ (Puterman, , 2014).
Let $(\mathcal{S},\mathcal{A},p,\mu,\gamma)$ be a Markov control process (MPC; Puterman, , 2014), i.e., an MDP without a reward function. Given an MPC, we define a family of MDPs $\mathcal{M}\triangleq\{(\mathcal{S},\mathcal{A},p,r,\mu,\gamma)\mid r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}\}$ that only differ in their reward function. We refer to any MDP $M\in\mathcal{M}$ (and its corresponding reward function $r$) as a task.
@@ 2.2 Generalized Policy Evaluation and Improvement
Generalized Policy Evaluation (GPE) and Generalized Policy Improvement (GPI) generalize the policy evaluation and improvement steps to the case where an agent has access to a set of policies. In particular, GPE and GPI are used, respectively, (i) to evaluate a policy on multiple tasks; and (ii) to construct a policy (appropriate for solving a particular novel task) by improving upon an existing set of policies.
Definition 2.1 (Barreto et al., (2020)).
GPE is the computation of the action-value function of a policy $\pi$, $q^{\pi}_{r}(s,a)$, on a set of tasks. Given a set of policies $\Pi$ and a reward function of an arbitrary task, $r$, GPI defines a policy, $\pi^{\prime}$, such that
$q^{\pi^{\prime}}_{r}(s,a)\geq\max_{\pi\in\Pi}q^{\pi}_{r}(s,a)\ \text{ for all }(s,a)\in\mathcal{S}\times\mathcal{A}.$ (1)
Based on this definition, for any reward function $r$, a GPI policy can be constructed based on a set of policies $\Pi$ as:
@@
Let $q^{\text{GPI}}_{r}(s,a)$ be the action-value function of $\pi^{\text{GPI}}$ under the reward function $r$. The GPI theorem (Barreto et al., , 2017) ensures that $\pi^{\text{GPI}}$ in Eq. (2) satisfies Def. 2.1; i.e., that $q^{\text{GPI}}_{r}(s,a)\geq\max_{\pi\in\Pi}q_{r}^{\pi}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. This implies that Eq. (2) can be used to define a policy guaranteed to perform at least as well as all other policies $\pi_{i}\in\Pi$ w.r.t. any given reward function, $r$. The GPI theorem can also be extended to the case $q^{\pi_{i}}$ is replaced with an approximation, $\tilde{q}^{\pi_{i}}$ (Barreto et al., , 2018).
@@ 2.3 GPE&GPI via Successor Features
Successor features (SFs) allows us to perform GPE&GPI efficiently (Barreto et al., , 2017). Assume that the reward functions of interest are linear w.r.t. reward features, $\bm{\mathbf{\phi}}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$. That is, a reward function $r_{\bm{\mathbf{w}}}$ can be expressed as $r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}$, where $\bm{\mathbf{w}}\in\mathbb{R}^{d}$ is a weight vector. Then, the SFs of a policy $\pi$ for a given state-action pair $(s,a)$, $\bm{\mathbf{\psi}}^{\pi}(s,a)\in\mathbb{R}^{d}$, is defined as
@@
where $\bm{\mathbf{\phi}}_{t}\triangleq\bm{\mathbf{\phi}}(S_{t},A_{t},S_{t+1})$. Notice that the definition of SFs corresponds to a form of action-value function, where the features $\bm{\mathbf{\phi}}_{t}$ play the role of rewards. Thus, SFs can be learned through any temporal-difference (TD) learning algorithm (e.g., Q-learning (Watkins, , 1989; Mnih et al., , 2015)). We write $\bm{\mathbf{\psi}}^{\pi}$ to denote the SF vector associated with $\pi$, where the expectation is with respect to the initial state distribution: $\bm{\mathbf{\psi}}^{\pi}\triangleq\mathbb{E}_{S_{0}\sim\mu}\left[\bm{\mathbf{\psi}}^{\pi}(S_{0},\pi(S_{0}))\right]$.
Given the SFs $\bm{\mathbf{\psi}}^{\pi}(s,a)$ of a policy $\pi$, it is possible to directly compute the action-value function $q^{\pi}_{\bm{\mathbf{w}}}(s,a)$ of $\pi$, under any linearly-expressible reward functions, $r_{\bm{\mathbf{w}}}$, as follows: $q_{\bm{\mathbf{w}}}^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\gamma^{i}r_{\bm{\mathbf{w}}}(S_{t+i},A_{t+i},S_{t+i+1})\mid S_{t}=s,A_{t}=a\right]=\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}.$ That is, given any set of reward weights, $\{\bm{\mathbf{w}}_{i}\}_{i=1}^{n}$, GPE can be performed efficiently via inner-products between the SFs and the reward weights: $q_{\bm{\mathbf{w}}_{i}}^{\pi}(s,a)=\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}_{i}$. Note that the value of a policy $\pi$ under any task $\bm{\mathbf{w}}$ can be expressed as $v^{\pi}_{\bm{\mathbf{w}}}=\bm{\mathbf{\psi}}^{\pi}\cdot\bm{\mathbf{w}}$.
Let $\Pi=\{\pi_{i}\}_{i=1}^{n}$ be a set of policies with corresponding SFs $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$. Based on the definition of GPI (Def. 2.1) and the reward decomposition $r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}$, a generalized policy, $\pi^{\text{GPI}}:\mathcal{S}\times\mathcal{W}\to\mathcal{A}$, can then be defined as follows:
@@ 2.4 GPI via the Option Keyboard
The Option Keyboard (OK; Barreto et al., , 2019, 2020) is a method that generalizes the GPI policy (Eq. (2)) by allowing the agent to follow a different weight vector at each time step. That is, the OK learns a meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ that outputs weights, $\bm{\mathbf{z}}\in\mathcal{Z}$, [^1] Without loss of generality, we let $\mathcal{Z}$ be the space of $d$-dimensional unit vectors, i.e., $\mathcal{Z}=\{\bm{\mathbf{z}}\in\mathbb{R}^{d}\mid||\bm{\mathbf{z}}||_{2}=1\}$. that are used when following the GPI policy $\pi^{\text{GPI}}(s,\bm{\mathbf{z}};\Pi)$ at each state $s$. Intuitively, given a set of policies $\Pi$, the policy $\omega$ modulates the behavior induced by $\pi^{\text{GPI}}$ via $\Pi$ by controlling the preferences $\omega(s)$ over features for each state $s$. In this case, the OK policy, $\pi^{\text{OK}}_{\omega}$, at each state $s$, is given by:
@@
Also note that $\pi_{\omega}^{\text{OK}}(s;\Pi)=\pi^{\text{GPI}}(s,\omega(s);\Pi)$. Intuitively, in the “keyboard” analogy, $\omega(s)$ (a “chord”) combines base policies $\Pi$ (“keys”) to generate more complex behaviors. A key property of the OK is that learning a policy over $\mathcal{Z}$ is often easier than learning one over the original action space $\mathcal{A}$, as solutions typically allow the same action $\bm{\mathbf{z}}$ to be repeated for multiple timesteps, similar to temporally extended options Sutton et al., (1999).
@@ 3 Optimal Behavior Basis
In what follows, we consider weight vectors that induce convex combinations of features; that is, $\mathcal{W}=\{\bm{\mathbf{w}}\mid\sum_{i}w_{i}=1$, $w_{i}\geq 0,\forall i\}$. This is a common practice, e.g., in the multi-objective RL literature (Hayes et al., , 2022), since the optimal policy of any MDP remains unchanged when its reward function is scaled by a positive constant, as in $\mathcal{W}$. Existing algorithms that optimally solve all tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ typically identify a set of policies, $\Pi_{k}=\{\pi_{i}\}_{i=1}^{n}$, such that their associated SF vectors, $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$, form a convex coverage set (CCS; Alegre et al., , 2022). A CCS is a set of SF vectors that allows the optimal value function $v^{*}_{\bm{\mathbf{w}}}=\bm{\mathbf{\psi}}^{\pi^{*}_{\bm{\mathbf{w}}}}\cdot\bm{\mathbf{w}}$ for any given task $\bm{\mathbf{w}}\in\mathcal{W}$ to be directly identified: [^2] A CCS, as defined in Eq. (7), may not be unique since tasks may allow for many optimal policies—each with a possibly different SF vector. In what follows, mentions of a CCS refer specifically to the minimal set satisfying Eq.(7), which is guaranteed to be unique (Roijers et al., , 2013).
@@
Unfortunately, methods that compute the complete CCS for $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ do not scale to complex tasks and become impractical as the number of reward features $d$ grows. Hence, it becomes critical to develop novel techniques capable of recovering all solutions induced by a CCS without incurring the cost of learning the corresponding full set of policies. This could be achieved, for instance, by methods capable of expressing all solutions in a CCS by combining policies in a behavior basis $\Pi_{k}$ smaller than the CCS; that is, $|\Pi_{k}|\leq|\mathrm{CCS}|$.
@@
Intuitively, the meta-policy $\omega(s,\bm{\mathbf{w}})$ enables the OK to decompose the optimal policy for a task $\bm{\mathbf{w}}$ by dynamically assigning state-dependent linear weights to the SFs of its various base policies. Note that if $\omega(s,\bm{\mathbf{w}})=\bm{\mathbf{w}}$ for all $s\in\mathcal{S}$, we recover GPI.
Given an arbitrary set of base policies $\Pi_{k}$, we define the space of all policies expressible by the OK, $\Pi^{\text{OK}}(\Pi_{k})$, and their associated SF vectors, $\Psi^{\text{OK}}(\Pi_{k})$, respectively, as
$\Pi^{\text{OK}}(\Pi_{k})$$\triangleq\{\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi_{k})\mid\bm{\mathbf{w}}\in\mathcal{W},\omega:\mathcal{S}\times\mathcal{W}\to\mathcal{Z}\},$
$\Psi^{\text{OK}}(\Pi_{k})$$\triangleq\{\bm{\mathbf{\psi}}^{\pi}\mid\pi\in\Pi^{\text{OK}}(\Pi_{k})\}.$ (9)
@@
As observed by Barreto et al., (2019, 2020), learning a meta-policy, $\omega$, is often easier than learning base policies over the original action space, $\mathcal{A}$. Consider, e.g., that if a task $\bm{\mathbf{w}}\in\mathcal{W}$ can be solved by switching between two base policies in $\Pi_{k}$, then an optimal meta-policy for $\bm{\mathbf{w}}$, $\omega(\cdot,\bm{\mathbf{w}})$, only needs to output two vectors $\bm{\mathbf{z}}$ (one for each base policy) for any given state. Thus, solving the goal above is bound to be more efficient than constructing a complete CCS since it requires learning fewer optimal base policies.

### v2

@@ 1 Introduction
Reinforcement learning (RL) methods have been successfully used to solve complex sequential decision-making problems (Silver+2017; Bellemare+2020). However, traditional RL algorithms typically require thousands or millions of interactions with the environment to learn a single policy for a single task. Multi-task RL methods address this limitation by enabling agents to quickly identify solutions for new tasks with minimal or no additional interactions. Such methods often achieve this by learning a set of specialized policies (or behavior basis) designed for specific tasks and subsequently combining them to more rapidly solve novel tasks.
@@
A powerful approach for combining policies to solve new tasks in a zero-shot manner leverages successor features (SFs) and generalized policy improvement (GPI) (Barreto+2018; Barreto+2020). GPI can combine base policies to solve a new task, producing a policy that is at least as good as any individual base policy. Importantly, however, GPI policies are not guaranteed to be optimal. Optimality can be ensured if tasks can be expressed as a linear combination of reward features using techniques that compute a convex coverage set (CCS) (Alegre+2022). A CCS is a set of policies that includes an optimal policy for any linear task. Unfortunately, the number of policies in a CCS often grows exponentially with the number of reward features (Roijers2016), making existing algorithms computationally expensive and difficult to scale to complex domains (Yang+2019; Alegre+2022; Alegre+2023).
@@
The option keyboard (OK) method improves upon GPI by producing policies that are at least as good—and often better. It achieves this through a learned meta-policy that dynamically combines base policies by assigning state-dependent linear weights to reward features (Barreto+2019). This allows OK to express a larger spectrum of behaviors than GPI and potentially solve tasks beyond linear rewards, enabling it to produce policies that GPI cannot represent. However, its performance critically depends on the choice of base policies available. Existing OK-based methods assume a predefined set of base policies or rely on expert knowledge to construct one, without addressing how to identify a good behavior basis (Barreto+2019; Carvalho+2023neurips). This raises a key question and open problem: is there an optimal set of base policies for the OK—an optimal behavior basis—that enables zero-shot identification of optimal solutions for any linear tasks?
@@
We solve this open problem by introducing Option Keyboard Basis (OKB), a novel method with strong formal guarantees that efficiently identifies an optimal behavior basis for the OK. OKB provably identifies a set of policies that allows the OK to optimally solve any linear task; i.e., it can express all policies in a CCS. We show that it significantly reduces the number of base policies required to ensure zero-shot optimality in new tasks. Furthermore, we prove that the set of policies it can express is strictly more expressive than a CCS, enabling particular classes of non-linear tasks to be solved optimally. We empirically evaluate our method in challenging high-dimensional RL problems and show that it consistently outperforms state-of-the-art GPI-based approaches. Importantly, we also observe that the performance gain over competing methods becomes more pronounced as the number of reward features increases.
@@ 2.1 Reinforcement Learning
An RL problem (Sutton&Barto2018) is typically modeled as a Markov decision process (MDP). An MDP is defined as a tuple $M\triangleq(\mathcal{S},\mathcal{A},p,r,\mu,\gamma)$, where $\mathcal{S}$ is a state space, $\mathcal{A}$ is an action space, $p(\cdot|s,a)$ describes the distribution over next states given that the agent executed action $a$ in state $s$, $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function, $\mu$ is an initial state distribution, and $\gamma\in[0,1)$ is a discounting factor. Let $S_{t}$, $A_{t}$, and $R_{t}\triangleq r(S_{t},A_{t},S_{t+1})$ be random variables corresponding to the state, action, and reward, respectively, at time step $t$. The goal of an RL agent is to learn a policy $\pi:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted sum of rewards (return), $G_{t}=\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}$. The action-value function of a policy $\pi$ is defined as $q^{\pi}(s,a)\triangleq\mathbb{E}_{\pi}[G_{t}|S_{t}=s,A_{t}=a]$, where $\mathbb{E}_{\pi}[\cdot]$ denotes the expectation over trajectories induced by $\pi$. Given $q^{\pi}$, one can define a greedy policy $\pi^{\prime}(s)\in\operatorname*{arg\,max}_{a}q^{\pi}(s,a)$. It is guaranteed that $q^{\pi^{\prime}}(s,a)\geq q^{\pi}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}$. The processes of computing $q^{\pi}$ and $\pi^{\prime}$ are known, respectively, as the policy evaluation and policy improvement steps. Under certain conditions, repeatedly executing the policy evaluating and improvement steps leads to an optimal policy $\pi^{*}(s)\in\operatorname*{arg\,max}\nolimits_{a}q^{*}(s,a)$ (Puterman2014). Let $(\mathcal{S},\mathcal{A},p,\mu,\gamma)$ be a Markov control process (MPC) (Puterman2014), i.e., an MDP without a reward function. Given an MPC, we define a family of MDPs $\mathcal{M}\triangleq\{(\mathcal{S},\mathcal{A},p,r,\mu,\gamma)\mid r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}\}$ that only differ in their reward function. We refer to any MDP $M\in\mathcal{M}$ (and its corresponding reward function $r$) as a task.
@@ 2.2 Generalized Policy Evaluation and Improvement
Generalized Policy Evaluation (GPE) and Generalized Policy Improvement (GPI) generalize the policy evaluation and improvement steps to the case where an agent has access to a set of policies. In particular, GPE and GPI are used, respectively, (i) to evaluate a policy on multiple tasks; and (ii) to construct a policy, capable of solving a particular novel task, by improving on an existing set of policies.
Definition 2.1 (Barreto+2020).
GPE is the computation of the action-value function of a policy $\pi$, $q^{\pi}_{r}(s,a)$, for a set of tasks. Moreover, given a set of policies $\Pi$ and a reward function of an arbitrary task, $r$, GPI defines a policy, $\pi^{\prime}$, such that $q^{\pi^{\prime}}_{r}(s,a)\geq\max_{\pi\in\Pi}q^{\pi}_{r}(s,a)\ \text{ for all }(s,a)\in\mathcal{S}\times\mathcal{A}.$
Based on the latter definition, for any reward function $r$, a GPI policy can be constructed based on a set of policies $\Pi$ as
@@
Let $q^{\text{GPI}}_{r}(s,a)$ be the action-value function of $\pi^{\text{GPI}}$ under the reward function $r$. The GPI theorem (Barreto+2017) ensures that $\pi^{\text{GPI}}$ in Eq. (1) satisfies Def. 2.1; i.e., that $q^{\text{GPI}}_{r}(s,a)\geq\max_{\pi\in\Pi}q_{r}^{\pi}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. This implies that Eq. (1) can be used to define a policy guaranteed to perform at least as well as all other policies $\pi_{i}\in\Pi$ w.r.t. any given reward function, $r$. The GPI theorem can also be extended to the case $q^{\pi_{i}}$ is replaced with an approximation, $\tilde{q}^{\pi_{i}}$ (Barreto+2018).
@@ 2.3 GPE&GPI via Successor Features
Successor features (SFs) allows us to perform GPE&GPI efficiently (Barreto+2017). Assume that the reward functions of interest are linear w.r.t. reward features, $\bm{\mathbf{\phi}}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$. That is, a reward function, $r_{\bm{\mathbf{w}}}$, can be expressed as $r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}$, where $\bm{\mathbf{w}}\in\mathbb{R}^{d}$ is a weight vector. Then, the SFs of a policy $\pi$ for a given state-action pair $(s,a)$, $\bm{\mathbf{\psi}}^{\pi}(s,a)\in\mathbb{R}^{d}$, are defined as
@@
where $\bm{\mathbf{\phi}}_{t}\triangleq\bm{\mathbf{\phi}}(S_{t},A_{t},S_{t+1})$. Notice that the definition of SFs corresponds to a form of action-value function, where the features $\bm{\mathbf{\phi}}_{t}$ play the role of rewards. Thus, SFs can be learned through any temporal-difference (TD) learning algorithm (e.g., Q-learning (Watkins1989; Mnih+2015) for learning $\bm{\mathbf{\psi}}^{\pi^{*}}$). We refer to $\bm{\mathbf{\psi}}^{\pi}\triangleq\mathbb{E}_{S_{0}\sim\mu}\left[\bm{\mathbf{\psi}}^{\pi}(S_{0},\pi(S_{0}))\right]$ as the SF vector associated with $\pi$, where the expectation is with respect to the initial state distribution.
Given the SFs $\bm{\mathbf{\psi}}^{\pi}(s,a)$ of a policy $\pi$, it is possible to directly compute the action-value function $q^{\pi}_{\bm{\mathbf{w}}}(s,a)$ of $\pi$, under any linearly-expressible reward functions, $r_{\bm{\mathbf{w}}}$, as follows: $q_{\bm{\mathbf{w}}}^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\gamma^{i}r_{\bm{\mathbf{w}}}(S_{t+i},A_{t+i},S_{t+i+1})\mid S_{t}=s,A_{t}=a\right]=\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}.$ That is, given any set of reward weights, $\{\bm{\mathbf{w}}_{i}\}_{i=1}^{n}$, GPE can be performed efficiently via inner-products between the SFs and the reward weights: $q_{\bm{\mathbf{w}}_{i}}^{\pi}(s,a)=\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}_{i}$. Note that the value of a policy $\pi$ under any task $\bm{\mathbf{w}}$ can be expressed as $v^{\pi}_{\bm{\mathbf{w}}}=\bm{\mathbf{\psi}}^{\pi}\cdot\bm{\mathbf{w}}$. Let $\Pi=\{\pi_{i}\}_{i=1}^{n}$ be a set of policies with corresponding SFs $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$. Based on the definition of GPI (Def. 2.1) and the reward decomposition $r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}$, a generalized policy, $\pi^{\text{GPI}}:\mathcal{S}\times\mathcal{W}\to\mathcal{A}$, can then be defined as follows:
@@ 2.4 GPI via the Option Keyboard
[…]
@@
Note that, by construction, $\pi_{\omega}^{\text{OK}}(s;\Pi)=\pi^{\text{GPI}}(s,\omega(s);\Pi)$. Intuitively, in the “keyboard” analogy, $\omega(s)$ (a “chord”) combines base policies in $\Pi$ (“keys”) to generate more complex behaviors. A key property of OK is that learning a policy over $\mathcal{Z}$ is often easier than over $\mathcal{A}$, as solutions typically involve repeating the same action $\bm{\mathbf{z}}$ across multiple timesteps, similar to temporally extended options.
@@ 3 Optimal Behavior Basis
Each element $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ (or, equivalently, its weight vector $\bm{\mathbf{w}}$) is called a task. In what follows, we consider weight vectors that induce convex combinations of features; that is, $\mathcal{W}=\{\bm{\mathbf{w}}\mid\sum_{i}w_{i}=1$, $w_{i}\geq 0,\forall i\}$. This is common practice, e.g., in the multi-objective RL literature (Hayes+2022). Our analyses naturally extend to linear combinations of features—in App. B, we show how linear-reward MDPs can be transformed into equivalent convex-reward MDPs (w.l.o.g.) so that our results and algorithms apply directly. Existing algorithms capable of optimally solving all tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ often identify a set of policies, $\Pi_{k}=\{\pi_{i}\}_{i=1}^{n}$, such that their associated SF vectors, $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$, form a convex coverage set (CCS) (Alegre+2022). A CCS is a set of SF vectors that allows the optimal value function, $v^{*}_{\bm{\mathbf{w}}}=\bm{\mathbf{\psi}}^{\pi^{*}_{\bm{\mathbf{w}}}}\cdot\bm{\mathbf{w}}$, for any given task $\bm{\mathbf{w}}\in\mathcal{W}$, to be directly identified: [^2] A CCS may not be unique since tasks can have multiple optimal policies with distinct SF vectors. In what follows, CCS refers specifically to the minimal set satisfying Eq. (6), which is uniquely defined (Roijers+2013).
@@
Unfortunately, methods that compute the complete CCS for $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ do not scale to complex tasks since its size can grow exponentially with dd—counting corner weights is equivalent to Vertex Enumeration (Roijers2016). Hence, it becomes critical to develop novel techniques capable of recovering all solutions induced by a CCS without incurring the cost of learning the corresponding full set of policies. This could be achieved, for instance, by methods capable of expressing all solutions in a CCS by combining policies in a behavior basis $\Pi_{k}$ smaller than the CCS; that is, $|\Pi_{k}|\leq|\mathrm{CCS}|$.
@@
The meta-policy $\omega(s,\bm{\mathbf{w}})$ enables the OK to decompose the optimal policy for a task $\bm{\mathbf{w}}$ by dynamically assigning state-dependent linear weights to the SFs of its various base policies. Note that if $\omega(s,\bm{\mathbf{w}})=\bm{\mathbf{w}}$ for all $s\in\mathcal{S}$, we recover GPI. Given an arbitrary set of base policies $\Pi_{k}$, we define the space of all policies expressible by the OK, $\Pi^{\text{OK}}(\Pi_{k})$, and their associated SF vectors, $\Psi^{\text{OK}}(\Pi_{k})$, respectively, as
$\Pi^{\text{OK}}(\Pi_{k})\triangleq\{\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi_{k})\mid\bm{\mathbf{w}}\in\mathcal{W},\omega:\mathcal{S}\times\mathcal{W}\to\mathcal{Z}\}\textrm{ and}\ \Psi^{\text{OK}}(\Pi_{k})\triangleq\{\bm{\mathbf{\psi}}^{\pi}\mid\pi\in\Pi^{\text{OK}}(\Pi_{k})\}.$
@@
As observed by Barreto+2019; Barreto+2020, learning a meta-policy, $\omega$, is often easier than learning base policies over the original action space, $\mathcal{A}$. Consider, e.g., that if a task $\bm{\mathbf{w}}\in\mathcal{W}$ can be solved by switching between two base policies in $\Pi_{k}$, then an optimal meta-policy for $\bm{\mathbf{w}}$, $\omega(\cdot,\bm{\mathbf{w}})$, only needs to output two vectors $\bm{\mathbf{z}}$ (one for each base policy) for any given state. Thus, solving the goal above is bound to be more efficient than constructing a complete CCS since it requires learning fewer optimal base policies.